OpenStudy (anonymous):

Patrick was sick and had to take 50 mg of antibiotics each day for 2 weeks. Each day he takes his medicine the amount in his body doubles. How much medicine will be in his body at the end of the 2 weeks?

OpenStudy (jack1):

day 1: 50 = a x n^0 (n=2, a=50) day 2: 100 = a x n^1 day 3: 200 = a x n^2 ... continue the series to get day 14: x = a x n^13

OpenStudy (jack1):

sorry, bad choice of terms day 14: y = a * n^13

OpenStudy (jack1):

so solve for y ...let us know if u dont understand something

OpenStudy (anonymous):

i dont understand how to solve for y

OpenStudy (jack1):

gotcha y = a * n^13 and: a = 50 n = 2 so \[y = a \times n ^{13}\]

OpenStudy (jack1):

sub in a = 50 and n = 2 into the above equation

OpenStudy (jack1):

\[ y = 50 \times 2^{13}\]

OpenStudy (jack1):

=...? @haiven ?

OpenStudy (anonymous):

oh okay i get it now (: